OpenStudy (babalooo):

URGENT!!! Just realized my answer was incorrect. if I have integral 1/(sqrt(1+v^2)dv = integral1/x dx where v = y/x I simplified the equation to arcsinhv = lnx +c With the initial condition at point (2,0), how do I find arbitrary constant and get y explicitly in terms of x?

OpenStudy (anonymous):

my bad. its sinh not sin. modifying answer.

OpenStudy (anonymous):

plug in (x, y) = (2, 0) and solve for C so arcsinh(0/2) = ln(2) + C C = - ln(2) and you can plug that C in your solution: you get: \[\sinh^{-1}\left( \frac{ y }{ x } \right) = \ln(x) - \ln(2)\]\[\frac{ y }{ x } = \sinh[\ln(x) - \ln(2)]\]\[y = xsinh[\ln(x) - \ln(2)]\]

OpenStudy (babalooo):

THANKS! LIFESAVER!

OpenStudy (anonymous):

happy to help ^_^